Why, in my finance class, when asked this question - You invest $1,000 for 3 years at 6% per annum compounding quarterly. What is the equivalent annual simple rate of interest? - am I told the answer is 6.521%, when my calculator shows 6.136%, when I preform (1195.62/1000)^(1*3)-1. Am I missing something when trying to work out interest? Any answers would be greatly appreciated. Thank you.
Found 2 solutions by jorel1380, MathTherapy:
Answer by jorel1380(3719)

The total interest accrued after 3 years of quarterly compounding would be:
(1 + .06/4)^(3*4)=(1.015)^12=1.1956181714615352515612900979004
1.1956181714615352515612900979004 - 1= 0.19561817% after 3 years
0.19561817/3=.065206, or an effective simple interest rate of 6.5206%. ☺☺☺☺

Answer by MathTherapy(10552)


Why, in my finance class, when asked this question - You invest $1,000 for 3 years at 6% per annum compounding quarterly. What is the equivalent annual simple rate of interest? - am I told the answer is 6.521%, when my calculator shows 6.136%, when I preform (1195.62/1000)^(1*3)-1. Am I missing something when trying to work out interest? Any answers would be greatly appreciated. Thank you.
The formula for future value, based on SIMPLE INTEREST is: A = P + PTR, where:
A = Future Value (1,195.62, in this case)
P = Principal (1,000, in this case)
T = Time, in years (3, in this case)
R = Annual interest RATE (Unknown, in this case)
Substitute these variables, solve for R, and see what you get!